# A reduced-order framework for temperature estimation in food freezing from optimally located sensors, including turbulent conjugate flow scenarios

### v1

@@ 3.1 Finite volume method
To compute a numerical solution, the system of equations (1)-(3), and (4) is discretized in space using the finite volume method (FVM) by decomposing the computational domain into $N_{v}$ non-overlapping cells $\Omega_{i}$. Integrating the PDEs over each cell and applying the divergence theorem yields:
@@
where $\phi$ stands for any state variable ($\boldsymbol{u},T,k,\omega$), $\Gamma$ is the corresponding diffusion coefficient, $\boldsymbol{f}$ stands for the corresponding forcing term, $S_{i}$ is the cell boundary, and $\boldsymbol{n}$ is the outward normal unit vector to $S_{i}$. We use a non-uniform structured mesh with a staggered grid approach that stores the cell-averaged value of scalars at the cell centroid and velocities at the cell faces.
The cell face reconstruction is based on a third-order weighted essentially non-oscillatory method (WENO3) [55] that uses two adaptive weights, $w_{0},w_{1}\geq 0$, with $w_{0}+w_{1}=1$, and two-point sub-stencils, $S_{i}^{0}(\phi_{i-1},\phi_{i})$ and $S_{i}^{1}(\phi_{i},\phi_{i+1})$, such that: $\phi_{S_{i}}=w_{0}S_{i}^{0}+w_{1}S_{i}^{1}$. Each sub-stencil approximates $\phi_{S_{i}}$ using a linear interpolation through the respective stencil points (cell centroids) in the non-uniform mesh. The weights are computed as a function of the state variable as:
@@
where $d_{1}=\frac{1}{3}$, $d_{2}=\frac{2}{3}$, and $\epsilon=10^{-6}$. Finally, we use the Lax–Friedrichs procedure to ensure a consistent monotone physical flux at the cell faces, and we use a second-order central difference method for the gradient reconstruction at cell faces.
@@
Velocity and pressure are coupled using the segregated pressure-correction method SIMPLEC, and the set of resulting linearized algebraic equations is solved by the Bi-Conjugated Stabilized Gradient method (Bi-CGSTAB) with the ILU preconditioner. The maximum relative difference of all state variables between two successive iterations is used as converge criteria. The solution converges when that error is less than $10^{-6}$.
@@ 3.2 Verification of the finite volume solver
    Freezing process of beef-meat inside a conventional freezer cabinet by turbulent natural convection of air, with a Rayleigh number of $Ra=6.81\times 10^{7}$ [21], and
@@
Figure 3 $a)$ shows, for P1, the numerical results for the temperature at two control points in the food domain ($A$ located at the geometrical center and $B$ located at a quarter height from the lower surface) comparing the numerical results obtained with the method described in Section 3 with previous results from the authors [10] and with available experimental temperature measurements. Concerning the results reported in [10], the present method uses higher order schemes for the cell-face reconstruction and for the time integration (WENO3 and BDF2-opt) and includes a $k$-$\omega$ SST model that significantly improves the accuracy of results, reducing the RMS error from 3.88 °C to 2.41 °C at $T_{A}$, and from 4.27 °C to 1.86 °C at $T_{B}$.
@@ 3.3 Forward numerical simulations
This Section describes the results of a representative parameter set defined as the mean values of the range for each boundary condition, i.e., $U_{in}=0.2$ m/s, $T_{C}=-22\ ^{\circ}\text{C}$, $T_{ext}=22\ ^{\circ}\text{C}$, and $h_{ext}=0.8\ \text{W}\text{m}^{-2}\text{K}^{-1}$. We discuss the main features of the direct numerical results, focusing, in particular, on the challenges related to the forced heat convection within a turbulent airflow and to the liquid-to-solid phase change process of the food.
@@
Figure 4 illustrates snapshots of the temperature, air velocity as $U=\sqrt{u_{1}^{2}+u_{2}^{2}}$, and turbulent kinetic energy fields at characteristic time instants: $t=0.001$ h (hours), $t=1.5$ h, $t=4.5$ h, and $t=8$ h. The results indicate that the inlet airflow causes rapid cooling of the air domain within the first minute, with some high-temperature stratifying isotherms in the upper region produced by natural convection effects. Contrary, the food cools down slowly due to its high thermal inertial caused by the latent heat during phase change. Moreover, both air velocity and turbulent kinetic energy fields indicate that at the very beginning the flow is more unsteady and chaotic, evidencing that the fluctuations and gradients of air velocity are notably higher. However, the flow stabilizes rapidly within the first hour and reaches a semi-steady state.
@@
To understand the different time scales of the fluid dynamics and heat transfer processes, Figure 5 depicts the variation in time of volume-average and maximum values of temperatures (air and food), of food liquid fraction ($F_{pc}$), and of time-averaged and instantaneous fluctuations of velocity, i.e. $U$ and $U^{\prime}=\sqrt{(2/3)k}$. Analyzing the figure one sees the abrupt drop in air temperatures. In contrast, the average and maximum temperature in the food illustrates a characteristic freezing curve where the effect of the latent heat in the temperature-slope is clearly defined in the freezing stage. Regarding the evolution of volume-average and maximum values the two components of the air velocity in a URANS model, the most dynamic patterns of $U$ and $U^{\prime}$ are observed at the first seconds of the freezing process, but then these values quickly stabilize until they reach a steady behavior over time. Moreover, the elevated value of $U_{max}$ is about four times higher than its volume-average value, demonstrating the presence of high-velocity gradients in the flow domain. Lastly, the relation between the instantaneous fluctuation and the time-average velocity, or turbulent intensity $I_{t}=U^{\prime}/U$, reaches volume-average and maximum values of 22% and 38%, respectively, indicating the convective-dominant nature of the problem. These points evidence the need for a suitable turbulent model that precisely computes the augmented diffusion of momentum and heat due to the turbulent mixing phenomenon.
@@
Figure 6 (left) shows snapshots of the temperature and liquid fraction fields within the food at characteristic time instants: $t=0.2$ h (hours), $t=1$ h, $t=3$ h, and $t=5$ h. As expected, the liquid fraction field evolution is directly correlated with the isotherms since it is a temperature-derived quantity. Another relevant aspect observed in the results is that the freezing center moves during the freezing process. Initially, the high-temperature isotherms are concentrated just below the geometric center, and during the phase change, they are rather concentrated close to the right surface of the food. This behavior occurs because the cold air flow slips along the upper wall and hits the food on the left, while the lower right corner is located precisely in a flow recirculation region, which weakens heat transfer.
To highlight the spatial differences in the dynamics of the food freezing process, Figure 6 (right) depicts the temporal evolution of temperature, liquid fraction, and volumetric heat capacity at two fixed points within the food domain. These points are conveniently selected as the freezing center ($P1$), and in the upper corner region ($P2$) where the freezing rate is expected to be highest. The difference in the path of temperature and liquid fraction evolution between the two points within the food shows the relevance of the locality in the phase change phenomenon. The center region experiences a more isothermal process than in the edges of the food, indicating a lower heat transfer rate in that area. Consequently, the maximum values of volumetric heat capacity related to the latent heat occur more slowly and at different periods of time in the freezing process.
@@ 4 Inverse problem
As schematically depicted in Figure 7, measuring a high dimensional temperature state is understood as applying the observer matrix $\boldsymbol{W}$ on a temperature state. The matrix columns, otherwise understood as the Riesz representers of the measurements [42], are indicatrix functions on the data pixels, i.e., for every thermo-camera pixel $\Omega_{i}\subset\Omega$ we have:
@@
Figure 7: The observation model as a linear operation on the target state. Local averages of the temperature are computed at each pixel of the thermo-camera data $\Omega_{i}\subset\Omega$.
@@ 4.2 Reduced-order modeling
This section discusses the generation of the reduced-order model $\boldsymbol{\Phi}$ using a parametrized singular value decomposition (SVD) [61] of a set of snapshots from the forward problem (named in certain contexts PCA – principal component analysis – or POD – proper orthogonal decomposition – in its generalization to arbitrary infinite-dimensional functional settings). The goal is to define a low-dimensional space that can sufficiently well approximate the physics of the phase change problem and the turbulent heat convection developed by the airflow. For this purpose, the first step consists of concatenating snapshots
@@
$\boldsymbol{A}=\boldsymbol{\Phi}\Sigma\boldsymbol{V}^{T}=\underbrace{\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ \vdots&\vdots&\ddots&\vdots\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\end{bmatrix}}_{\boldsymbol{\Phi}\ (\mathcal{N}\times n)}\underbrace{\begin{bmatrix}\sigma_{1}&0&\cdots&0\\ 0&\sigma_{2}&\cdots&0\\ \vdots&\vdots&\ddots&0\end{bmatrix}}_{\Sigma\ (n\times n)}\underbrace{\begin{bmatrix}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&\cdots&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}\\ {\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&\cdots&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}\\ \vdots&\vdots&\ddots&\vdots&\vdots\end{bmatrix}^{T}}_{\boldsymbol{V}^{T}\ (n\times k)}$ (11)

### v2

@@ 3 Direct problem
This section is devoted to describing the numerical approach we adopt to approximate solutions of the differential equations (4). We select a finite volume scheme, introduced next in detail, along with the temporal discretization of the problem. Next, we test our forward solver with a beef freezing set-up, benchmarking the numerical outcomes with available localized experimental measurements.
@@ 3.1 Finite volume method
To compute a numerical solution, the system of equations (1), (2),(4) is discretized in space using the finite volume method (FVM) by decomposing the computational domain into $N_{v}$ non-overlapping cells $\Omega_{i}$. Integrating the PDEs over each cell and applying the divergence theorem yields:
@@
where $\phi$ stands for any state variable ($\boldsymbol{u},T,k,\omega$), $\Gamma$ is the corresponding diffusion coefficient, $\boldsymbol{f}$ stands for the corresponding forcing term, $S_{i}$ is the cell boundary, and $\boldsymbol{n}$ is the outward normal unit vector to $S_{i}$. We use a structured anisotropic Cartesian mesh with a staggered grid approach that stores the cell-averaged value of scalars at the cell centroid and velocities at the cell faces.
We selected the Weighted Essentially Non-Oscillatory (WENO) scheme because of its advantages over other commonly used higher-order schemes such as the Total Variation Diminishing (TVD) family, from both numerical and implementation standpoints. The WENO scheme achieves arbitrarily high-order accuracy in smooth regions by combining nonlinear weights and lower-order stencils, maintaining stable and non-oscillatory solutions. Furthermore, using a monotonic flux function (Godunov or Lax-Friedrichs, for example), spurious oscillations (undershoots/overshoots) are eliminated, ensuring monotonicity-preserving methods [77, 78]. This issue is particularly important when sharp gradients in temperature or velocity occur. Also, implementing the WENO method is straightforward in anisotropic Cartesian structured meshes, a key property for mechanical science applications. The cell face reconstruction used is based on a third-order WENO method (WENO3) [79] that combines two adaptive weights, $w_{0},w_{1}\geq 0$, with $w_{0}+w_{1}=1$, and two-point sub-stencils, $S_{i}^{0}(\phi_{i-1},\phi_{i})$ and $S_{i}^{1}(\phi_{i},\phi_{i+1})$, such that: $\phi_{S_{i}}=w_{0}S_{i}^{0}+w_{1}S_{i}^{1}$. For incompressible turbulent flows such as those studied here, including dissipation is a must, and the key feature is the correct dissipative structure of the method and its ability to represent physically admissible solutions, which WENO3 fulfills [80, 81, 82, 83, 84], as discussed in the direct model validation section. For the WENO3 approach, each sub-stencil approximates $\phi_{S_{i}}$ using a linear interpolation through the respective stencil points (cell centroids) in the anisotropic mesh. The weights are computed as a function of the state variable as:
@@
where $d_{0}=\frac{1}{3}$, $d_{1}=\frac{2}{3}$, and $\epsilon=10^{-6}$. Finally, we use the Lax–Friedrichs procedure to ensure a consistent monotone physical flux at the cell faces, and we use a second-order central difference method for the gradient reconstruction at cell faces.
@@
Velocity and pressure are coupled using the segregated pressure-correction method SIMPLEC, and the set of resulting linearized algebraic equations is solved by the Bi-Conjugated Stabilized Gradient method (Bi-CGSTAB) with the ILU preconditioner. The SIMPLEC algorithm offers a significant advantage over other segregation methods, such as SIMPLE or PISO, particularly when addressing highly nonlinear problems. This advantage stems from its lower computational cost per outer iteration, as only one inner iteration is needed, which enhances the overall nonlinear convergence and the coupling between the unknown fields. The problem being solved involves the coupling of turbulent viscosity and buoyancy forces in the linear momentum equation, as well as temperature-dependent properties of food in the energy equation. This study effectively used a fixed-point iterative strategy to linearize the problem, making it straightforward to implement within the SIMPLEC segregation method. The maximum relative difference of all state variables between two successive iterations is used as converge criteria. The solution converges when that error is less than $10^{-6}$.
Regarding the boundary conditions, Dirichlet-type ones are imposed on the finite volume nodes, and Neumann boundary conditions are defined using first-order finite differences considering non-uniform grids. Robin conditions are also computed with first-order finite differences. Near-wall grid refinement using a tangent hyperbolic function was essential to deal with the imposition of turbulent boundary conditions. Grid refinement allows us to define the first cell height (which dimensionless is $y^{+}$) close to the recommended value for URANS turbulent models, $y^{+}=\mu_{t}\delta x_{n}/\nu\approx 1$.
@@ 3.2 Verification of the finite volume solver
    Freezing process of beef inside a conventional freezer cabinet by turbulent natural convection of air, with a Rayleigh number of $Ra=6.81\times 10^{7}$ [21], and
@@
For P2, Fig. 3$(b)$ compares the numerical results for the considered configuration with the experimental profiles of vertical air velocity at the mid-height of the cavity $(Y=0.5)$ and the Nusselt number at the hot wall $(X=0)$. Despite the thin boundary layer, the results agree with the experimental data, obtaining RMS error values of 0.05 ms-1 and 18.3 for the velocity and Nusselt numbers, respectively. From these and the previous results, we conclude that the considered Finite Volume Method can accurately describe the turbulent convection of air and the food freezing process.
@@ 3.3 Forward numerical simulations
In this Section we discuss the main features of the forward numerical results related to the forced heat convection within a turbulent airflow and to the phase change process of the food. We use a representative parameter set defined as the mean values of the range for each boundary condition, i.e., $U_{in}=0.2$ ms-1, $T_{C}=-22\ ^{\circ}\text{C}$, $T_{ext}=22\ ^{\circ}\text{C}$, and $h_{ext}=0.8\ \text{W}\text{m}^{-2}\text{K}^{-1}$.
To define the number of volume cells and the time step size required to ensure the independence of solutions regarding discretization, we conduct a convergence study by comparing velocity and temperature profiles along control lines. We define three different meshes composed of $18,410$ (M1 mesh), $22,802$ (M2 mesh), and $29,070$ (M3 mesh) cells. In this work, we use a wall-concentrated anisotropic Cartesian mesh defined to capture velocity and temperature gradients and accurately capture turbulent quantities. The meshes used were constructed by dividing the domain into eighteen subdomains to control the size of the control volumes separately. The minimum and maximum wall-normal lengths dimensions used to construct each mesh are defined as $h_{min,i}$ and $h_{max,i}$, being $i$, a counter that refers to the specific mesh. In specific, $h_{min,1}=4.0\cdot 10^{-3}$ and $h_{max,1}=8.5\cdot 10^{-2}$, $h_{min,2}=7.2\cdot 10^{-3}$ and $h_{max,2}=1.8\cdot 10^{-3}$, and $h_{min,3}=1.3\cdot 10^{-4}$ and $h_{max,3}=5.1\cdot 10^{-3}$. Concerning the discretization, we compare three different time step sizes of $\Delta t_{1}=0.03$ s, $\Delta t_{2}=0.05$ s, and $\Delta t_{3}=0.08$ s, and we evaluate the effect of this discretization parameter in the numerical results.
In Fig. 4(b), we compare the dimensionless vertical profiles of velocity and temperature at the freezer cabinet mid-plane at $t=5$ h (hours), using $U_{ref}=0.5$ ms-1. When comparing the root mean square error (RMS) at fixed points between each successive mesh size and time step, we found that the results changed on average by 2.1% from $M1$ to $M2$, and by 1.3% from $M2$ to $M3$. Regarding the time step, the results varied by 1.4% from $\Delta t_{1}$ to $\Delta t_{2}$, and by 0.5% from $\Delta t_{2}$ to $\Delta t_{3}$. Given the low variation of the results with the discretization parameters, we selected the mesh $M3$ and the time step $\Delta t_{2}$ to conduct the numerical test through the work.
Figure 4: (a) Mesh refinement of the selected mesh $M3$ with a zoom region near the food. (b) Discretization study: comparison of velocity and temperature profiles between three mesh sizes (upper) and three time steps (lower).
@@
Fig. 5 illustrates snapshots of the temperature, air velocity as $U=\sqrt{u_{1}^{2}+u_{2}^{2}}$, and turbulent kinetic energy fields at characteristic time instants: $t=0.001$ h, $t=1.5$ h, $t=4.5$ h, and $t=8$ h. The results indicate that the inlet airflow causes rapid cooling of the air domain within the first minute, with some high-temperature stratifying isotherms in the upper region produced by natural convection effects. Contrary, the food cools down slowly due to its high thermal inertial caused by the latent heat during phase change. Moreover, both air velocity and turbulent kinetic energy fields indicate that at the very beginning the flow is more unsteady and chaotic, evidencing that the fluctuations and gradients are notably higher. However, the flow stabilizes rapidly within the first hour and reaches a state of periodic oscillations of velocity and temperature that are unnoticeable since they occur on a small time scale.
@@
To understand the different time scales of the fluid dynamics and heat transfer processes, Fig. 6 depicts the time evolution of volume-average and maximum values of temperatures (air and food), of food liquid fraction ($F_{pc}$), and of time-averaged and instantaneous fluctuations of velocity (i.e. $U$ and $U^{\prime}=\sqrt{(2/3)k}$), revealing the abrupt drop in air temperatures. In contrast, the average and maximum temperature in the food illustrates a characteristic freezing curve where the effect of the latent heat on the temperature-slope is clearly defined in the freezing stage. Regarding the evolution of volume-average and maximum values the two components of the air velocity in a URANS model, the most dynamic patterns of $U$ and $U^{\prime}$ are observed at the first seconds of the freezing process, but then these values quickly stabilize until they reach a steady behavior over time. Moreover, the elevated value of $U_{max}$ is about four times higher than its volume-average value, demonstrating the presence of high-velocity gradients in the flow domain. Lastly, the relation between the instantaneous fluctuation and the time-average velocity, or turbulent intensity $I_{t}=U^{\prime}/U$, reaches volume-average and maximum values of 22% and 38%, respectively, indicating the convective-dominant nature of the problem. These points evidence the need for a suitable turbulent model that accurately computes the increased momentum and heat diffusion due to the turbulent mixing phenomenon.
@@
Fig. 7 (left) shows snapshots of the temperature and liquid fraction fields within the food at $t=0.2$ h, $t=1$ h, $t=3$ h, and $t=5$ h. As expected, the liquid fraction field evolution is directly correlated with the isotherms since it is a temperature-derived quantity. Another relevant aspect observed in the results is that the freezing center moves during the freezing process. Initially, the high-temperature isotherms are concentrated just below the geometric center, and during the phase change, they are rather concentrated close to the right surface of the food. This behavior occurs because the cold airflow slips along the upper wall and hits the food on the left, while the lower right corner is located precisely in a flow recirculation region, which weakens heat transfer.
To highlight the spatial differences in the dynamics of the food freezing process, Fig. 7 (right) depicts the temporal evolution of temperature, liquid fraction, and volumetric heat capacity at two fixed points within the food domain. These points are conveniently selected as the freezing center ($P1$), and in the upper corner region ($P2$) where the freezing rate is expected to be highest. The difference in the path of temperature and liquid fraction evolution between the two points within the food shows the relevance of the locality in the phase change phenomenon. The interior of the food experiences a more isothermal process than the periphery, indicating a lower heat transfer rate. Consequently, the maximum values of volumetric heat capacity related to the latent heat occur more slowly and at different periods of time in the freezing process.
@@ 4 Inverse problem
As schematically depicted in the right bottom corner of Fig. 8, measuring a high dimensional temperature state is understood as applying the observer matrix $\boldsymbol{W}$ on a temperature state. The matrix columns, otherwise understood as the Riesz representers of the measurements [56], are indicatrix functions on the data pixels, i.e., for every thermo-camera pixel $\Omega_{i}\subset\Omega$ we have:
@@ 4.2 Reduced-order modeling
This section discusses the generation of the reduced-order model $\boldsymbol{\Phi}$ using a parametrized singular value decomposition (SVD) of a set of snapshots from the forward problem (named in certain contexts PCA – principal component analysis – or POD – proper orthogonal decomposition – in its generalization to arbitrary infinite-dimensional functional settings). The goal is to define a low-dimensional space that can sufficiently well approximate the physics of the phase change problem and the turbulent heat convection developed by the airflow. For this purpose, the first step consists of concatenating snapshots
@@
$\boldsymbol{A}=\boldsymbol{\Phi}\Sigma\boldsymbol{V}^{T}=\underbrace{\begin{bmatrix}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ \vdots&\vdots&\ddots&\vdots\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}&\cdots&{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\blacksquare}\end{bmatrix}}_{\boldsymbol{\Phi}\ (\mathcal{N}\times n)}\underbrace{\begin{bmatrix}\sigma_{1}&0&\cdots&0\\ 0&\sigma_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\sigma_{n}\end{bmatrix}}_{\Sigma\ (n\times n)}\underbrace{\begin{bmatrix}{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&\cdots&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}\\ {\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&\cdots&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}&{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\blacksquare}\\ \vdots&\vdots&\ddots&\vdots&\vdots\end{bmatrix}^{T}}_{\boldsymbol{V}^{T}\ (n\times k)},$ (11)